# CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations

### v2

@@ Appendix B Statistical Analysis
We present two summary tables for output prediction tasks (Table 8) and input prediction tasks (Table 9). The results show that across a total of $357$ measurements, only 9 cases have $\sigma>1.0$ (i.e., $0.01$ on a scale from 00 to $1$), and $57$ have $\sigma>0.5$, indicating that most scores exhibit low variance under repeated sampling.
Furthermore, our benchmark includes $1279$ problems across $8$ experiment settings on $17$ models for output prediction tasks, and $800$ problems across $5$ experiment settings on $17$ models for input prediction tasks. Among $241,944$ combinations, $231,726$ have identical correctness labels across all three candidates, performing an overall stability rate of $\mathbf{95.78\%}$.
@@
Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
@@
Qwen2.5-Coder-32B-Instruct $0.4602$ $0.2604$ $0.4714$ $0.5905$ $0.1768$ $0.858$ $0.5803$ $0.1705$ $0.2041$ $0.7873$ $0.4249$ $0.2952$ 00 $0.3409$ $0.3118$ $0.6889$
@@
Model Series Model Name VAN ALL MDC MPS MHC
@@
Appendix C Prompts
Our prompt instructions for LLMs follow the format designed by Crux, which is also adopted by LCB, utilizing a two-shot prompt for direct inference and a single-shot prompt for CoT inference to ensure models can follow the instruction and output format effectively. For LRMs, internal reasoning is inherently required before responding. Therefore, applying CoT prompts may lead to redundant reasoning and an unfair comparison with standard LLMs under CoT prompting. Considering the fairness, we use direct inference prompts for RM experiments in §4.3.2.
Unlike LCB, we rename all example function names to f to avoid prompting models to capture the NL cues from the example. We add the instruction “Do NOT output any extra information” to prevent external output, particularly from Gemini and Claude models, to ensure all models have a fair output environment. We also include “Ensure the provided expression syntax is correct” to enforce consistent syntax. However, some responses from models like QwQ-32B, DeepSeek-R1 still generate syntactically incorrect outputs such as [ANSWER] f("a") == [/ANSWER] instead of [ANSWER] f("a") == "" [/ANSWER]. For those situations, we identify them as “Syntax Errors” and incorrect.
@@ Appendix C Prompts
C.1 Output Prediction (Direct)
@@ C.2 Output Prediction (CoT)
C.3 Input Prediction (Direct)
@@ C.4 Input Prediction (CoT)
Appendix D Supplementary Experimental Results
D.1 Impact on Output Prediction
@@
In §4.2, we evaluate the robustness of LLMs on output prediction tasks under direct inference. Table 1 reports the absolute Pass@1 degradation across all perturbations. We use absolute values to highlight the overall impact of each perturbation and conduct cross-perturbation comparisons. For completeness, we also include the relative Pass@1 degradation in Table 10, reflecting the percentage drop from each model’s vanilla performance.
@@
Table 10: Relative Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in direct inference ($N=3$). [To Absolute Version]
Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $57.29$ $53.24$ $-1.09\%$ $-18.17\%$ $-3.13\%$ $-1.83\%$ $-19.05\%$ $-10.33\%$ $-27.85\%$ $-31.90\%$ $-32.58\%$ $-40.39\%$ $-23.20\%$ $-27.97\%$ $-11.56\%$ $-44.31\%$
GPT-4o $71.33$ $64.51$ $-3.80\%$ $-14.89\%$ $-4.73\%$ $-7.12\%$ $-12.91\%$ $-4.96\%$ $-15.01\%$ $-28.37\%$ $-13.96\%$ $-29.13\%$ $-17.23\%$ $-24.27\%$ $-6.37\%$ $-26.65\%$
Claude Series Claude-3.5-Haiku-20241022 $57.42$ $58.25$ $-6.89\%$ $-16.01\%$ $-3.12\%$ $+0.60\%$ $-12.70\%$ $-7.77\%$ $-24.02\%$ $-36.92\%$ $-13.79\%$ $-10.04\%$ $-11.25\%$ $-8.00\%$ $-27.50\%$ $-53.29\%$
Claude-3.5-Sonnet-20241022 $71.46$ $73.83$ $-2.33\%$ $-22.71\%$ $-1.28\%$ $+1.60\%$ $-9.62\%$ $-6.69\%$ $-14.75\%$ $-34.12\%$ $-8.10\%$ $-9.61\%$ $-8.63\%$ $-10.65\%$ $-14.40\%$ $-43.45\%$
Gemini Series Gemini-1.5-Flash $56.21$ $44.26$ $-6.30\%$ $-0.94\%$ $-2.08\%$ $-2.36\%$ $-8.15\%$ $-0.79\%$ $-18.38\%$ $-12.74\%$ $-14.75\%$ $-32.23\%$ $-28.47\%$ $-40.41\%$ $-7.34\%$ $-22.17\%$
Gemini-2.0-Flash $65.04$ $65.97$ $-3.59\%$ $-27.43\%$ $-3.20\%$ $-1.16\%$ $-23.83\%$ $-13.29\%$ $-33.25\%$ $-42.09\%$ $-26.78\%$ $-51.90\%$ $-33.12\%$ $-50.32\%$ $-8.26\%$ $-19.51\%$
Gemini-1.5-Pro-002 $67.37$ $56.16$ $-4.33\%$ $-6.07\%$ $-3.71\%$ $-8.43\%$ $-8.35\%$ $-6.82\%$ $-19.91\%$ $-23.17\%$ $-23.01\%$ $-32.96\%$ $-21.71\%$ $-32.84\%$ $-10.02\%$ $-32.47\%$
DeepSeek DeepSeek-V3 $67.87$ $67.78$ $-3.93\%$ $-24.13\%$ $-1.96\%$ $-5.54\%$ $-6.81\%$ $-8.83\%$ $-12.89\%$ $-35.63\%$ $-16.57\%$ $-41.38\%$ $-10.13\%$ $-34.19\%$ $-10.74\%$ $-29.67\%$
LLaMA Series LLaMA-3.1-8B-Instruct $35.96$ $34.66$ $+0.23\%$ $-12.45\%$ $-0.93\%$ $-4.82\%$ $-8.69\%$ $-13.65\%$ $-23.75\%$ $-19.68\%$ $-22.48\%$ $-28.51\%$ $-17.84\%$ $-23.09\%$ $-16.69\%$ $-39.16\%$
LLaMA-3.1-70B-Instruct $56.08$ $44.89$ $-4.23\%$ $-3.10\%$ $-0.97\%$ $-3.88\%$ $-9.58\%$ $-5.43\%$ $-19.61\%$ $-17.21\%$ $-19.24\%$ $-27.91\%$ $-26.67\%$ $-38.76\%$ $-8.40\%$ $-32.25\%$
LLaMA-3.1-405B-Instruct $63.54$ $50.66$ $-8.46\%$ $-7.69\%$ $-0.98\%$ $-8.79\%$ $-9.11\%$ $-5.63\%$ $-19.67\%$ $-16.90\%$ $-6.56\%$ $-14.70\%$ $-11.28\%$ $-19.92\%$ $-6.89\%$ $-25.69\%$
LLaMA-3.3-70B-Instruct $59.92$ $48.50$ $-3.69\%$ $-2.87\%$ $-2.09\%$ $-7.03\%$ $-9.18\%$ $-11.76\%$ $-17.32\%$ $-20.09\%$ $-12.45\%$ $-19.80\%$ $-13.49\%$ $-25.39\%$ $-6.82\%$ $-20.95\%$
Qwen Series Qwen2.5-7B-Instruct $43.29$ $41.41$ $-7.03\%$ $-17.48\%$ $-2.98\%$ $-7.73\%$ $-29.93\%$ $-9.08\%$ $-37.92\%$ $-30.92\%$ $-58.04\%$ $-38.32\%$ $-45.24\%$ $-19.83\%$ $-26.95\%$ $-55.63\%$
Qwen2.5-14B-Instruct $47.79$ $49.48$ $-3.84\%$ $-19.97\%$ $-1.39\%$ $-2.81\%$ $-23.45\%$ $-6.89\%$ $-39.84\%$ $-29.96\%$ $-34.87\%$ $-41.35\%$ $-32.52\%$ $-33.90\%$ $-9.33\%$ $-22.22\%$
Qwen2.5-32B-Instruct $60.04$ $59.64$ $-2.85\%$ $-18.09\%$ $-4.30\%$ $-0.82\%$ $-9.65\%$ $-2.33\%$ $-19.78\%$ $-34.89\%$ $-18.04\%$ $-32.67\%$ $-23.66\%$ $-23.57\%$ $-13.12\%$ $-30.81\%$
Qwen2.5-72B-Instruct $60.08$ $54.91$ $-1.73\%$ $-15.46\%$ $-3.12\%$ $+0.76\%$ $-17.82\%$ $-3.68\%$ $-23.30\%$ $-25.22\%$ $-16.99\%$ $-12.67\%$ $-24.06\%$ $-26.24\%$ $-16.02\%$ $-40.56\%$
Qwen2.5-Coder-32B-Instruct $66.96$ $56.65$ $-3.73\%$ $-14.86\%$ $-1.43\%$ $-3.56\%$ $-16.05\%$ $-4.91\%$ $-22.34\%$ $-27.89\%$ $-23.21\%$ $-36.61\%$ $-16.55\%$ $-30.34\%$ $-10.45\%$ $-21.74\%$
Average $-3.98\%$ $-14.25\%$ $-2.44\%$ $-3.70\%$ $-13.82\%$ $-7.23\%$ $-22.92\%$ $-27.51\%$ $-21.26\%$ $-29.42\%$ $-21.47\%$ $-27.63\%$ $-12.40\%$ $-32.97\%$
$\mathbf{-7.83}\%$ $\mathbf{-2.91}\%$ $\mathbf{-11.35}\%$ $\mathbf{-24.64}\%$ $\mathbf{-24.32}\%$ $\mathbf{-23.78}\%$ $\mathbf{-20.11}\%$
D.2 Robustness Across Model Scales
In §4.2.3, we compare the robustness against different model scales. To fairly aggregate model robustness across different perturbation types, we compute a weighted average of the relative Pass@1 degradation under four representative perturbations: ALL (structural), MDC and MPS (misleading messages), and MHC (hints). We select ALL as the representative structural perturbation because of the compounding effects of REN, RTF, and GBC captured in §4.2.2, which reflect the most challenging structural setting. Since MDC and MPS stem from the same perturbation strategy and often target overlapping AST node types, we assign them lower weights ($1/8$ each) to avoid over-representing similar effects. Overall, we define a weighted robustness metric with weights of $1/2$, $1/8$, $1/8$, and $1/4$ for ALL, MDC, MPS, and MHC, respectively, and summarize the results in Table 2. For completeness, we also report the results under uniform weighting ($1/4$ for each) in Table 11, which shows similar trends and confirms the stability of our findings in §4.2.3, as well as their absolute degradation versions in Table 12 and 13, respectively.
@@ D.2 Robustness Across Model Scales
Table 11: Relative Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/4 ALL, 1/4 MDC, 1/4 MPS, and 1/4 MHC. [To Absolute Version]
Model Series Model Name VAN ALL ($1/4$) MDC ($1/4$) MPS ($1/4$) MHC ($1/4$) Average
GPT-4 Omni GPT-4o-Mini 55.77 $-29.37\%$ $-35.51\%$ $-24.99\%$ $-23.83\%$ $-28.42\%$
GPT-4o 68.78 $-20.01\%$ $-19.64\%$ $-19.87\%$ $-13.96\%$ $-18.37\%$
Claude Series Claude-3.5-Haiku-20241022 57.73 $-28.85\%$ $-12.38\%$ $-10.03\%$ $-37.16\%$ $-22.11\%$
Claude-3.5-Sonnet-20241022 72.35 $-22.01\%$ $-8.67\%$ $-9.39\%$ $-25.28\%$ $-16.34\%$
Gemini Series Gemini-1.5-Flash 51.73 $-16.27\%$ $-21.30\%$ $-32.94\%$ $-12.89\%$ $-20.85\%$
Gemini-2.0-Flash 65.39 $-36.56\%$ $-36.19\%$ $-39.56\%$ $-12.48\%$ $-31.20\%$
Gemini-1.5-Pro-002 63.17 $-21.13\%$ $-26.73\%$ $-25.88\%$ $-18.43\%$ $-23.04\%$
DeepSeek DeepSeek-V3 67.84 $-21.41\%$ $-25.86\%$ $-19.14\%$ $-17.83\%$ $-21.06\%$
LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-22.23\%$ $-24.74\%$ $-19.81\%$ $-25.10\%$ $-22.97\%$
LLaMA-3.1-70B-Instruct 51.89 $-18.71\%$ $-22.49\%$ $-31.20\%$ $-17.33\%$ $-22.43\%$
LLaMA-3.1-405B-Instruct 58.72 $-18.63\%$ $-9.61\%$ $-14.51\%$ $-13.93\%$ $-14.17\%$
LLaMA-3.3-70B-Instruct 55.64 $-18.35\%$ $-15.20\%$ $-17.95\%$ $-12.11\%$ $-15.90\%$
Qwen Series Qwen2.5-7B-Instruct 42.59 $-35.30\%$ $-50.65\%$ $-35.72\%$ $-37.69\%$ $-39.84\%$
Qwen2.5-14B-Instruct 48.42 $-36.14\%$ $-37.30\%$ $-33.04\%$ $-14.16\%$ $-30.16\%$
Qwen2.5-32B-Instruct 59.89 $-25.44\%$ $-23.52\%$ $-23.63\%$ $-19.74\%$ $-23.08\%$
Qwen2.5-72B-Instruct 58.14 $-24.02\%$ $-15.37\%$ $-24.88\%$ $-25.21\%$ $-22.37\%$
Qwen2.5-Coder-32B-Instruct 63.10 $-24.42\%$ $-28.23\%$ $-21.72\%$ $-14.68\%$ $-22.26\%$
Table 12: Absolute Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/2 ALL, 1/8 MDC, 1/8 MPS, and 1/4 MHC. [To Relative Version]
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Average
GPT-4 Omni GPT-4o-Mini 55.77 $-16.34$ $-19.73$ $-13.89$ $-12.98$ $-15.62$
GPT-4o 68.78 $-13.55$ $-13.27$ $-13.55$ $-9.28$ $-12.45$
Claude Series Claude-3.5-Haiku-20241022 57.73 $-16.68$ $-7.14$ $-5.79$ $-21.50$ $-15.33$
Claude-3.5-Sonnet-20241022 72.35 $-16.03$ $-6.28$ $-6.80$ $-18.45$ $-14.26$
Gemini Series Gemini-1.5-Flash 51.73 $-8.57$ $-10.53$ $-16.71$ $-6.25$ $-9.26$
Gemini-2.0-Flash 65.39 $-23.92$ $-23.72$ $-25.91$ $-8.18$ $-20.21$
Gemini-1.5-Pro-002 63.17 $-13.27$ $-16.63$ $-16.05$ $-11.05$ $-13.48$
DeepSeek DeepSeek-V3 67.84 $-14.52$ $-17.54$ $-12.98$ $-12.09$ $-14.10$
LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-7.90$ $-8.76$ $-7.01$ $-8.84$ $-8.13$
LLaMA-3.1-70B-Instruct 51.89 $-9.77$ $-11.44$ $-15.87$ $-8.37$ $-10.39$
LLaMA-3.1-405B-Instruct 58.72 $-11.02$ $-5.39$ $-8.26$ $-7.61$ $-9.12$
LLaMA-3.3-70B-Instruct 55.64 $-10.14$ $-8.26$ $-9.67$ $-6.36$ $-8.90$
Qwen Series Qwen2.5-7B-Instruct 42.59 $-15.06$ $-21.66$ $-15.32$ $-15.92$ $-16.14$
Qwen2.5-14B-Instruct 48.42 $-17.46$ $-18.09$ $-16.00$ $-6.91$ $-14.72$
Qwen2.5-32B-Instruct 59.89 $-15.22$ $-14.07$ $-14.15$ $-11.81$ $-14.09$
Qwen2.5-72B-Instruct 58.14 $-13.94$ $-8.99$ $-14.44$ $-14.36$ $-13.49$
Qwen2.5-Coder-32B-Instruct 63.10 $-15.27$ $-17.49$ $-13.37$ $-8.99$ $-13.74$
Table 13: Absolute Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/4 ALL, 1/4 MDC, 1/4 MPS, and 1/4 MHC. [To Relative Version]
Model Series Model Name VAN ALL ($1/4$) MDC ($1/4$) MPS ($1/4$) MHC ($1/4$) Average
GPT-4 Omni GPT-4o-Mini 55.77 $-16.34$ $-19.73$ $-13.89$ $-12.98$ $-15.73$
GPT-4o 68.78 $-13.55$ $-13.27$ $-13.55$ $-9.28$ $-12.41$
Claude Series Claude-3.5-Haiku-20241022 57.73 $-16.68$ $-7.14$ $-5.79$ $-21.50$ $-12.78$
Claude-3.5-Sonnet-20241022 72.35 $-16.03$ $-6.28$ $-6.80$ $-18.45$ $-11.89$
Gemini Series Gemini-1.5-Flash 51.73 $-8.57$ $-10.53$ $-16.71$ $-6.25$ $-10.52$
Gemini-2.0-Flash 65.39 $-23.92$ $-23.72$ $-25.91$ $-8.18$ $-20.43$
Gemini-1.5-Pro-002 63.17 $-13.27$ $-16.63$ $-16.05$ $-11.05$ $-14.25$
DeepSeek DeepSeek-V3 67.84 $-14.52$ $-17.54$ $-12.98$ $-12.09$ $-14.28$
LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-7.90$ $-8.76$ $-7.01$ $-8.84$ $-8.12$
LLaMA-3.1-70B-Instruct 51.89 $-9.77$ $-11.44$ $-15.87$ $-8.37$ $-11.36$
LLaMA-3.1-405B-Instruct 58.72 $-11.02$ $-5.39$ $-8.26$ $-7.61$ $-8.07$
LLaMA-3.3-70B-Instruct 55.64 $-10.14$ $-8.26$ $-9.67$ $-6.36$ $-8.61$
Qwen Series Qwen2.5-7B-Instruct 42.59 $-15.06$ $-21.66$ $-15.32$ $-15.92$ $-16.99$
Qwen2.5-14B-Instruct 48.42 $-17.46$ $-18.09$ $-16.00$ $-6.91$ $-14.61$
Qwen2.5-32B-Instruct 59.89 $-15.22$ $-14.07$ $-14.15$ $-11.81$ $-13.81$
Qwen2.5-72B-Instruct 58.14 $-13.94$ $-8.99$ $-14.44$ $-14.36$ $-12.93$
Qwen2.5-Coder-32B-Instruct 63.10 $-15.27$ $-17.49$ $-13.37$ $-8.99$ $-13.78$
D.3 Effect of CoT Reasoning
In §4.3.1, we analyze the effect of CoT prompting on model robustness using relative degradation aggregated across all perturbations (Table 3). We adopt the relative metric to mitigate the large variance between vanilla performances in direct and CoT inference modes, providing a fair comparison of performance degradation. For completeness, we also report the absolute degradation in Table 14, as well as the relative and absolute degradation on both Crux and LCB in Tables 15 and 16, respectively.
Table 14: Comparison of absolute Pass@1 degradation aggregated over Crux and LCB for output prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Relative Version]
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average

### v3

@@ Appendix C Statistical Analysis
We present two summary tables for output prediction tasks (Table 11) and input prediction tasks (Table 12). The results show that across a total of $357$ measurements, only 9 cases have $\sigma>1.0$ (i.e., $0.01$ on a scale from 0 to $1$), and $57$ have $\sigma>0.5$, indicating that most scores exhibit low variance under repeated sampling.
Furthermore, our benchmark includes $1279$ problems across $8$ experiment settings on $17$ models for output prediction tasks, and $800$ problems across $5$ experiment settings on $17$ models for input prediction tasks. Among $241,944$ combinations, $231,726$ have identical correctness labels across all three candidates, yielding an overall stability rate of $\mathbf{95.78\%}$.
@@
Model Series Model Name VAN REN RTF GBC PSC-ALL MCC MPS MHC
@@
Qwen2.5-Coder-32B-Instruct $0.4602$ $0.2604$ $0.4714$ $0.5905$ $0.1768$ $0.858$ $0.5803$ $0.1705$ $0.2041$ $0.7873$ $0.4249$ $0.2952$ 0 $0.3409$ $0.3118$ $0.6889$
@@
Model Series Model Name VAN PSC-ALL MCC MPS MHC
@@
Appendix D Prompts
Our prompt instructions for LLMs follow the format designed by Crux, which is also adopted by LCB, utilizing a two-shot prompt for direct inference and a single-shot prompt for CoT inference to ensure models can follow the instruction and output format effectively. For LRMs, internal reasoning is inherently required before responding. Therefore, applying CoT prompts may lead to redundant reasoning and an unfair comparison with standard LLMs under CoT prompting. Considering the fairness, we use direct inference prompts for LRM experiments in §3.4.
Unlike LCB, we rename all example function names to f to avoid prompting models to capture the NL cues from the example. We add the instruction “Do NOT output any extra information” to prevent external output, particularly from Gemini and Claude models, to ensure all models have a fair output environment. We also include “Ensure the provided expression syntax is correct” to enforce consistent syntax. However, some responses from models like QwQ-32B, DeepSeek-R1, still generate syntactically incorrect outputs such as [ANSWER] f("a") == [/ANSWER] instead of [ANSWER] f("a") == "" [/ANSWER]. For those situations, we identify them as “Syntax Errors” and incorrect.
@@ Appendix D Prompts
D.1 Output Prediction (Direct)
@@ D.2 Output Prediction (CoT)
D.3 Input Prediction (Direct)
@@ D.4 Input Prediction (CoT)
Appendix E Supplementary Experimental Results
E.1 Additional Performance Degradation Tables for Output Prediction (Direct Inference)
@@
In §3.2, we evaluate the robustness of LLMs on output prediction tasks under direct inference, using the relative Pass@1 degradation in Table 1. For completeness, we also include the absolute Pass@1 degradation in Table 13, reflecting the percentage drop from each model’s vanilla performance.
@@
Table 13: Absolute Pass@1 degradation of LLMs under all perturbation types for output prediction tasks on Crux and LCB using direct inference. [Relative Version] for the relative degradation.
Model Series Model Name VAN PSC-ALL MCC MPS MHC Average
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $57.3$ $53.2$ 16.0-16.0 17.0-17.0 18.7-18.7 21.5-21.5 13.3-13.3 14.9-14.9 6.6-6.6 23.6-23.6 15.7-15.7
GPT-4o $71.3$ $64.5$ 10.7-10.7 18.3-18.3 10.0-10.0 18.8-18.8 12.3-12.3 15.7-15.7 4.5-4.5 17.2-17.2 12.4-12.4
Claude Series Claude-3.5-Haiku-20241022 $57.4$ $58.2$ 13.8-13.8 21.5-21.5 7.9-7.9 5.8-5.8 6.5-6.5 4.7-4.7 15.8-15.8 31.0-31.0 12.8-12.8
Claude-3.5-Sonnet-20241022 $71.5$ $73.8$ 10.5-10.5 25.2-25.2 5.8-5.8 7.1-7.1 6.2-6.2 7.9-7.9 10.3-10.3 32.1-32.1 11.9-11.9
Gemini Series Gemini-1.5-Flash $56.2$ $44.3$ 10.3-10.3 5.6-5.6 8.3-8.3 14.3-14.3 16.0-16.0 17.9-17.9 4.1-4.1 9.8-9.8 10.5-10.5
Gemini-2.0-Flash $65.0$ $66.0$ 21.6-21.6 27.8-27.8 17.4-17.4 34.2-34.2 21.5-21.5 33.2-33.2 5.4-5.4 12.9-12.9 20.4-20.4
Gemini-1.5-Pro-002 $67.4$ $56.2$ 13.4-13.4 13.0-13.0 15.5-15.5 18.5-18.5 14.6-14.6 18.4-18.4 6.7-6.7 18.2-18.2 14.2-14.2
DeepSeek DeepSeek-V3 $67.9$ $67.8$ 8.7-8.7 24.1-24.1 11.2-11.2 28.0-28.0 6.9-6.9 23.2-23.2 7.3-7.3 20.1-20.1 14.3-14.3
LLaMA Series LLaMA-3.1-8B-Instruct $36.0$ $34.7$ 8.5-8.5 6.8-6.8 8.1-8.1 9.9-9.9 6.4-6.4 8.0-8.0 6.0-6.0 13.6-13.6 8.1-8.1
LLaMA-3.1-70B-Instruct $56.1$ $44.9$ 11.0-11.0 7.7-7.7 10.8-10.8 12.5-12.5 15.0-15.0 17.4-17.4 4.7-4.7 14.5-14.5 11.4-11.4
LLaMA-3.1-405B-Instruct $63.5$ $50.7$ 12.5-12.5 8.6-8.6 4.2-4.2 7.4-7.4 7.2-7.2 10.1-10.1 4.4-4.4 13.0-13.0 8.1-8.1
LLaMA-3.3-70B-Instruct $59.9$ $48.5$ 10.4-10.4 9.7-9.7 7.5-7.5 9.6-9.6 8.1-8.1 12.3-12.3 4.1-4.1 10.2-10.2 8.6-8.6
Qwen Series Qwen2.5-7B-Instruct $43.3$ $41.4$ 16.4-16.4 12.8-12.8 25.1-25.1 15.9-15.9 19.6-19.6 8.2-8.2 11.7-11.7 23.0-23.0 17.0-17.0
Qwen2.5-14B-Instruct $47.8$ $49.5$ 19.0-19.0 14.8-14.8 16.7-16.7 20.5-20.5 15.5-15.5 16.8-16.8 4.5-4.5 11.0-11.0 14.6-14.6
Qwen2.5-32B-Instruct $60.0$ $59.6$ 11.9-11.9 20.8-20.8 10.8-10.8 19.5-19.5 14.2-14.2 14.1-14.1 7.9-7.9 18.4-18.4 13.8-13.8
Qwen2.5-72B-Instruct $60.1$ $54.9$ 14.0-14.0 13.8-13.8 10.2-10.2 7.0-7.0 14.5-14.5 14.4-14.4 9.6-9.6 22.3-22.3 12.9-12.9
Qwen2.5-Coder-32B-Instruct $67.0$ $56.6$ 15.0-15.0 15.8-15.8 15.5-15.5 20.7-20.7 11.1-11.1 17.2-17.2 7.0-7.0 12.3-12.3 13.8-13.8
Average 13.2-13.2 15.5-15.5 12.0-12.0 16.0-16.0 12.3-12.3 15.0-15.0 7.1-7.1 17.8-17.8
$\mathbf{-14.0}$ $\mathbf{-13.5}$ $\mathbf{-13.3}$ $\mathbf{-11.1}$ $\mathbf{-13.0}$
E.2 Additional Performance Degradation Tables for Output Prediction (CoT Inference)
In §3.3, we analyze the effect of CoT prompting on model robustness using relative degradation aggregated across all perturbations (Table 2). We adopt the relative metric to mitigate the large variance between vanilla performances in direct and CoT inference modes, providing a fair comparison of performance degradation. For completeness, we report the absolute degradation in Table 14, as well as the relative and absolute degradation on both Crux and LCB in Tables 15 and 16, respectively.
@@ E.2 Additional Performance Degradation Tables for Output Prediction (CoT Inference)
Table 14: Comparison of absolute Pass@1 degradation aggregated over Crux and LCB for output prediction in different reasoning mode. See [Relative Version] for the relative degradation.
Model Series Model Name VAN PSC-ALL MCC MPS MHC Average